OpenStudy (anonymous):

The vote count for candidate A was 7,310. The vote count for candidate B was 14,190. Only two candidates were running. What percent of voters voted for candidate A?

OpenStudy (anonymous):

Assuming 7310 + 14190 represents all the voters, then the percent for candidate A is: 7310/(7310+14190) x 100% = 34%

OpenStudy (campbell_st):

the total number of voters was 7310 + 14190 = 25000 candiate A = (7310/25000) x 100

OpenStudy (anonymous):

Thank you!! :)